820) A game consists of tossing 3 coins in which it costs $0.10 to play, with
  a reward of $1.00 by tossing all three heads.  What is the cost to
  play 24 games?  How much money do you expect to receive?  ans:2
Answer #1

Ans:

P(all three heads)=1/8

P(other than all three heads)=7/8

Cost to play a single game=0.10

Cost to play 24 games=24*0.10=2.4 dollars

expected amount of money you receive in 24 games=24*(1*(1/8)-0*(7/8))=24*0.125=3 dollars